OpenStudy (anonymous):

jim walked from his house to office at his normal speed. if he walks 1 kph faster, he will reach his office 1 hour and 40 mins earlier. if he walks 1kph slower, he will be late by 3 hours and 20 minutes. how far his is office from his home?

OpenStudy (anonymous):

from where r u copying these questions

OpenStudy (anonymous):

this is an old question i had which no one answered yet. :D

OpenStudy (anonymous):

r t = d Solve the following for r, t and d:\[\left\{r t=d,(r+1)\left( t-\left(1+\frac{40}{60}\right)\right) \text{ = } d,(r-1)\left(t+\left(3+\frac{20}{60}\right)\right)=d\right\}\]\[\left\{r=3\text{ kph},t=\frac{20}{3}=6\frac{2}{3}\text{ hours}, d=20\text{ km}\right\} \]